A person purchased a ​$198741 home 10 years ago by paying ​20% down and signing a​ 30-year mortgage at 11.4​% compounded monthly. Interest rates have dropped and the owner wants to refinance the unpaid balance by signing a new 20​-year mortgage at compounded 6.6% monthly. How much interest will refinancing​ save?
Money Saved: $

Please Help! Thank you :)

Found 3 solutions by ikleyn, MathTherapy, math_tutor2020:
Answer by ikleyn(52781)

.

It is written to inform you that at this forum  we do teach  Math.


At this forum,  we do not teach  Finance and do not consult in  Finance.


From your side,  it requires some intellectual skills to distinct,  which problem
relates to  Math and can be solved at this forum,
and which one relates to  Finance and can not be solved here.

So,  it requires some  IQ  level to make a right decision.


Have a nice day / nigh / morning / afternoon / tea time / evening  ( ! )



Answer by MathTherapy(10552)


A person purchased a ​$198741 home 10 years ago by paying ​20% down and signing a​ 30-year mortgage at 11.4​% compounded monthly. Interest rates have dropped and the owner wants to refinance the unpaid balance by signing a new 20​-year mortgage at compounded 6.6% monthly. How much interest will refinancing​ save?
Money Saved: $

Please Help! Thank you :)
Amount of mortgage loan: .8(198,741) = $158,992.80

Each monthly payment for 1st 10 years (to date): $1,562.37

Using the formula for the balance on the mortgage loan, we find the unpaid balance after 10 years of monthly 
payments (120 periods) to be: $147,456.55 

Over the next 20 years, and at a 6.6% annual interest rate, each monthly payment to pay off balance ($147,456.55), is $1,108.09

INITIAL payoff amount for 30 years: $562,453.20 (360 * 1,562.37)

INITIAL 30-year interest amount: $562,453.20 - 158,992.80 = $403,460.40 

Total amount paid over 10 years (120 months): 120(1,562.37) = $187,484.40

Initial principal/mortgage-loan: $158,992.80
Principal/mortgage-loan balance after 10 years (120 months): $147,456.55
Amount applied to initial principal/mortgage-loan in 10 years: $158,992.80 - 147,456.55 = 11,536.25
Amount applied to mortgage interest in 10 years: $187,484.40 - 11,536.25 = $175,948.15

Unpaid interest at the 10-year mark: $403,460.40 - 175,948.15 = $227,512.25

Total amount to be paid in 240 monthly periods, or in 20 years, at the NEW rate: 240(1,108.09) = $265,941.60
Interest to be paid from total paid in 240 months, or 20 years: $265,941.60 - 147,456.55 = $118,485.05

Interest saved: $227,512.25 - 118,485.05 = $109,027.20

Answer by math_tutor2020(3817)


I'm using the formulas found on this page
https://www.mtgprofessor.com/formulas.htm

The home value is $198,741.
20% of it is paid off through the downpayment, so 80% is remaining
0.80*198741 = 158,992.80

This is the loan amount L
L = 158,992.80

For the original mortgage at 11.4%, the monthly rate c is
c = r/12 = 0.114/12 = 0.0095
and the number of months is n = 360 because 30*12 = 360.

So,
L = 158,992.80
c = 0.0095
n = 360
Plug those values into the first formula mentioned in that link above. You should get a monthly payment of P = 1562.37

Plug those values into the second formula along with lowercase p = 120 (since 10 years = 120 months) to find that B = 147,456.55
Unfortunately that link uses uppercase P to represent payment and lowercase p to denote the number of months paid into; the letter choice may seem a bit confusing.

Anyways, we get B = 147,456.55 which is the unpaid balance at year 10. This is the starting point, and we'll have two branching options

----------------------------------------------------

starting balance = 147,456.55

Mortgage A (interest rate = 11.4%)
monthly payment = 1562.37
total amount paid back = (240 months)*(1562.37 per month) = 374,968.80
m = Total interest paid
m = 374,968.80 - 147,456.55
m = 227,512.25

Mortgage B (interest rate = 6.6%)
monthly payment = 1108.09
total amount paid back = (240 months)*(1108.09 per month) = 265,941.60
n = Total interest paid
n = 265,941.60 - 147,456.55
n = 118,485.05

m-n = amount of interest saved by going with mortgage B
m-n = 227,512.25 - 118,485.05
m-n = 109,027.20


Answer: $109,027.20